if i have a right triangle where 26 is the hypotenuse(C) and bottom leg is 21 (C). how do i find A, the side leg?
Answer by stanbon(75887)

if i have a right triangle where 26 is the hypotenuse(C)
and bottom leg is 21 (C).
how do i find A, the side leg?
-------
Pythagoras:
21^2 + A^2 = 26^2
---
A^2 = 26^2 - 21^2
----
A^2 = 235
---
A = 15.33
================
Cheers,
Stan H.